# VGA: Vision GUI Assistant - Minimizing Hallucinations through Image-Centric Fine-Tuning

### v2

@@ Abstract
Recent advances in Large Vision-Language Models (LVLMs) have significantly improve performance in image comprehension tasks, such as formatted charts and rich-content images. Yet, Graphical User Interface (GUI) pose a greater challenge due to their structured format and detailed textual information. Existing LVLMs often overly depend on internal knowledge and neglect image content, resulting in hallucinations and incorrect responses in GUI comprehension. To address these issues, we introduce VGA, a fine-tuned model designed for comprehensive GUI understanding. Our model aims to enhance the interpretation of visual data of GUI and reduce hallucinations. We first construct a Vision Question Answering (VQA) dataset of 63.8k high-quality examples with our propose Referent Method, which ensures the model’s responses are highly depend on visual content within the image. We then design a two-stage fine-tuning method called Foundation and Advanced Comprehension (FAC), to enhance both the model’s ability to extract information from image content and alignment with human intent. Experiments show that our approach enhances the model’s ability to extract information from images and achieves state-of-the-art results in GUI understanding tasks. Our dataset and fine-tuning script will be released soon.
@@
Figure 1: Data generation Method
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have recently emerged as a powerful approach for various multimodal tasks. These models acquire textual knowledge through pre-training and develop image understanding abilities during instruction-tuning Cao et al. (2023); Liu et al. (2023). LVLMs can effectively process both visual and linguistic information by using a visual-language projector, which maps different types of data into a shared latent space. For instance, LLaVA Liu et al. (2024a) and InstructBLIP Dai et al. (2023) have demonstrated strong zero-shot capabilities in tasks like image captioning, visual reasoning, and complex conversations. UniChart Masry et al. (2023), on the other hand, has demonstrated the powerful capabilities of LVLMs in the understanding of formatted charts.
With the proliferation of mobile applications, the importance of Graphical User Interfaces (GUIs), which serve as a critical bridge between end users and applications, has increasingly garnered scholarly attention. GUIs, characterized by their structured layouts, rich graphical and textual content, and the inclusion of human operational logic Banerjee et al. (2013), present a complex challenge: Can the success of LVLMs be applied to the GUI domain?
Existing GUI comprehension method has transferred from code-centric base to visual base. While code-centric base approaches offer a foundational outline Chen et al. (2020b), they fail to fully capture the essence of GUI content, making visual-based methods more advantageous Yu et al. (2020)Bajammal et al. (2022). Compared to code-centric methods, visual-based approaches can convey the actual user interface interactions and mirror the user’s direct experience with the GUI. For instance, LabelDroid Chen et al. (2020a) utilizes deep learning to predict labels for image-based buttons from a variety of commercial apps available on Google Play. Similarly, TANGO Cooper et al. (2021) employs custom computer vision (CV) and text retrieval techniques to analyze visual and textual information on mobile screens.
However, traditional visual GUI comprehension focuses on identifying GUI components and fails to comprehend the graphical and textual information, layout, and interaction context within the interface. True comprehension demands that the model interpret the GUI in a manner akin to human perception. Existing LVLMs demonstrate significant advantages in this domain as they can understand both formatted charts and rich textual content while leveraging natural language aligned with human interpretations Zhang et al. (2024). Nevertheless, directly applying conventional LVLMs to GUI comprehension tasks often results in hallucinations, where the model over-relies on its pre-trained knowledge and neglects the actual visual content, leading to inaccurate or irrelevant responses Shahgir et al. (2024); Zhang et al. (2023).
@@
  • Performance enhancements on real-world GUI tasks. As shown in experiment  3, we apply our VGA to a GUI comprehension bench, yielding promising results.
@@ 2.2 Large Vision-Language Models
Besides, using transformers to map visual information into a semantic space has ensured the robust extension of LLMs into multimodal applications Yin et al. (2023). This enables models to process vision information, exemplified by multimodal recognition models like CLIP Radford et al. (2021), which align semantic information from images with text in a unified vector space. Advanced methodologies by Sun et al. (2023) and Fang et al. (2023) have further refined these capabilities.
@@
3 GUI Dataset Construction
@@
Existing general GUI dataset cannot be directly use for LVLM training. To address this gap, we propose a novel approach to generating high-quality GUI comprehension dataset that meets the urgent demand of LVLM training.
3.1 Existing General Dataset
The Rico Deka et al. (2017) dataset is a huge GUI dataset created to support research in mobile app design and development, including areas such as GUI design, interaction. It consist of 66k unique GUI screens and 3M elements from 27 categories, over 9.3k applications. Each GUI comprise a screenshot and an augmented Android view hierarchy that capture all of the elements comprising a GUI, their properties, and relationships between them.
However, the rich data provided by Rico cannot be directly utilized to train LVLMs as the text data in Rico do not align with human perception in the same manner as VQA datasets. This necessitates the transformation of this data into a format compatible with VQA dataset.
3.2 Propesed GUI VQA Dataset Generation
Inspired by the works like LLaVA Chen et al. (2024) and ShareGPT4V Chen et al. (2023), we use state-of-the-art models like GPT-4OpenAI (2024) to generate training data. More details about this approach can be found in Appendix A.1
In constructing high quality VQA datasets for GUI comprehension tasks, multiple quality dimensions must be considered:
Interaction: The ability to understand user interactions with GUI elements is crucial for the model to comprehend GUI functionality.
Positional: Accurate description of element locations within the interface is crucial, as relative position imply connections between GUI elements, such as buttons next to text boxes indicating related actions.
Visual Detail: Accurate representation of visual information, such as shape and color, is crucial. This allows the model to distinguish between elements and focus on their design attributes within the GUI.
Overall Function: Comprehensive overall functional descriptions enable the model to grasp the intended purpose and usage patterns of the interface, facilitating more accurate and effective comprehension.
The Android view hierarchy in the Rico dataset provides detailed descriptions of GUI elements, allowing us to obtain precise positional data and interaction details. Therefore, we use GPT-4 (text-only) to extract this textual information and GPT-4o to extract visual information as shown in Figure 1. Our approaches are as follows:
@@ 3.2 Propesed GUI VQA Dataset Generation
  • Data Generation by GPT-4 (text-only): Inspired by the method of LLaVA Liu et al. (2024b), We leverage the advanced capabilities of GPT-4 models (gpt-4-0125-preview and gpt-4-1106-preview) to generate VQA data. We employ these models to parse Android View hierarchy, extract precise element locations and user interaction information. Our approach ensures a accurate description of GUIs, including positional relationships and user interaction.
  • Data Generation by GPT-4o (vision-model): Inspired by ShareGPT4 Chen et al. (2023), we employ GPT-4o to directly process images and generate QA pairs based on the visual content. we capture the design patterns between different GUI elements.
Example
[Uncaptioned image] [Uncaptioned image]
User How can I give a like to this video? How can I login?
Find login section…
Enter your credentials…
@@
4 GUI Comprehension Task Design
In this section, we discuss how hallucinations arise in GUI understanding tasks and introduce our dataset design, explaining how it addresses this issue. The design of our dataset is related to our fine-tuning method. For more details, please see Section  5.
4.1 Hallucination in GUI Comprehension
The primary cause of hallucinations in LVLMs can be attributed to their dependence on response patterns learned by traditional Large Language Models (LLMs). LLMs are typically trained on massive corpora of pure text data, where they learn to generate responses based on textual context and patterns. When a visual module is introduced, these learned textual response patterns can adversely influence the model’s behavior during multimodal tasks, as shown in Table 2.
In a GUI, the structured layout conveys relationships between image and text elements, such as a number beneath a heart indicating likes. Misunderstanding these positional relationships can cause incorrect model responses. Moreover, action-indicating text often guides user behavior in GUI design, but the attention method might cause the model to over focus on similar words in the question. For instance, the model might suggest clicking a "login" button when you want to login in a page even if it doesn’t exist.
Referent Method For hallucination Reduction: To address the previously mentioned issue, we propose the referent method. GUI design often uses layout, shape, and color to distinguish elements, providing critical visual details to guide user interactions. By explicitly incorporating coordinates, shapes, colors, and relative positions between elements in the construction of our dataset, we aim to enhance the model’s focus on the image content when generating answers. Thereby reducing the chances of hallucination. During the design of our dataset, we ensured that most responses involving GUI elements include at least one of the following referents, aligning the element information with the image content:
4.2 Foundation Task For Knowledge Injection
The Foundation Task dataset is an instruction-following dataset that fix the command and response formats, ensuring that only the image content and response information vary dynamically. This dataset enhance model to identity the linking between specific data characteristics with visual attributes and understand the functions provided by the GUI. Our foundation task can be divided as follow:
  • Simple Description: This task involves providing a basic description of the GUI’s layout, identifying the function and placement of components. This foundational task facilitates deeper analytical tasks, as all GUI-related tasks rely on a clear understanding of the overall layout.
  • Overall Function: This task requires the model to grasp the individual functions of each GUI element based on overall function of the entire interface. Identical buttons might serve different purposes depending on their specific context within a GUI. This enables the model to predict the design intent and functionality of the entire interface, resulting in more contextually aware and precise responses.
4.3 Advanced Task For Human Alignment
Building upon the capabilities developed in the foundation tasks, the advanced comprehension tasks are designed to simulate real-world human interactions and complex dialogues within GUI comprehension tasks. This tasks are characterized by more flexible and contextually rich questions and responses compared to the foundational tasks. While the basic capabilities required are similar to those in the Foundation task, the advanced comprehension tasks emphasize the application and deeper understanding of these foundational skills. Our advanced comprehension tasks can be divided into two categories, distinguished by the generating models and its "referent":
  • GPT-4 (Text-Only) Generated Tasks: This category consists of conversations generated by GPT-4 (text-only), based on pre-processed Android View Hierarchy data. Leveraging the detailed information available in the Android View Hierarchy, this approach generates a variety of QA pairs that accurately reflect the textual and positional data and interactions.
  • GPT-4o (Image-Based) Generated Tasks: This category comprises dialogues generated by GPT-4o (image-based). Given that vision models often struggle with capturing precise positional information of elements, the question and response in this dataset focus on relative positions, shapes, and colors of elements.
5 Foundation and Advanced Comprehension
@@
Foundation Stage: In this stage, we fine-tune the model with the Foundation task dataset, which covers essential GUI comprehension capabilities. The fixed instruction and response format (detailed in sec 4.2) guarantees that only the images and response content vary for each sample. This approach makes response generation focusing more on image content, and builds foundational skills for GUI comprehension.
Advanced Comprehensive Stage: In this stage, we introduce multi-turn dialogues and complex questions from the advanced comprehension dataset, where each sample has a flexible question and answer format aligned with human intent.. In this phase, model learns human questioning patterns and utilizes foundational skills acquired from the first stage to generate responses. By employing the Referent Method (detailed in sec 4.1), our approach ensures that the model’s answers include more specific image details, thereby reducing hallucinations and improving the accuracy of the responses.
@@
Task Progression and Reinforcement: We organize tasks in a sequence from simple to complex, ensuring that the model progressively builds the necessary foundational skills before training on more advanced tasks. In complex tasks, we start with some related simple tasks to reinforce foundational knowledge, followed by the real response which demands deeper analysis and synthesis of the information gathered from these foundational tasks.
@@ 6.2 Main Result
FAC method analysis: As shown in Figure 2, the orange and purple curves represent the original loss values for models with and without the foundation task stage, respectively. The red and blue curves, fitted with a degree-3 polynomial, illustrate the loss trends. Model trained with foundation task exhibit more stable convergence and lower loss value. Figure 3 shows that mixing datasets during training leads to unstable convergence. In Appendix A.3, we discuses the influence of the foundation stage on model convergence. In Appendix A.4, we conducted ablation studies to evaluate the individual contributions of the two stages to overall model performance. In Appendix A.5, we analyzed the contribution of image tokens when generating answers to demonstrate how FAC enhances the model’s attention to image content.
@@ Acknowledgements
The authors would like to thank the anonymous reviewers for their helpful comments. This research was supported by Bytedance/TikTok, we would like to extend our gratitude to Bytedance/TikTok for their assistance and support.

### v3

@@ Abstract
Large Vision-Language Models (VLMs) have already been applied to the understanding of Graphical User Interfaces (GUIs) and have achieved notable results. However, existing VLMs often overly rely on internal text-based knowledge while neglecting visual inputs. This imbalance may lead models to produce answers that do not align with the visual content in GUI comprehension tasks. Such inaccuracies are termed as ‘hallucinations’ where models generate incorrect or illogical responses upon visual verification against GUI elements. These errors result in misinterpretations and diminish the model’s practical utility in applied settings. To address these issues, we introduce VGA, a fine-tuned model designed for comprehensive GUI understanding. Our model aims to balance attention image and text to enhance interpretation and reduce hallucinations. We construct a Vision Question Answering (VQA) dataset of 63.8k high-quality examples with our propose Referent Method, focusing on response with visual content of images. We then design a two-stage fine-tuning method to enhance both the model’s accuracy to extract information from image content and alignment with human intent. Experiments show that our approach enhances the model’s ability to extract information from images and achieves state-of-the-art results in GUI understanding tasks. Our dataset and fine-tuning script are available at https://github.com/Linziyang1999/VGA-visual-GUI-assistant
@@
Hallucination in GUI comprehension
[Uncaptioned image] [Uncaptioned image]
User How can I give a like to this video? How can I login?
Find login section…
Enter your credentials…
@@ 1 Introduction
Large Vision-Language Models (VLMs) have recently emerged as a powerful approach for various multimodal tasks. These models acquire textual knowledge through pre-training and develop image understanding abilities during instruction-tuning Cao et al. (2023); Liu et al. (2023). VLMs can effectively process both visual and linguistic information by using a visual-language projector, which maps different types of data into a shared latent space. For instance, LLaVA Liu et al. (2024a) and InstructBLIP Dai et al. (2023) have demonstrated strong zero-shot capabilities in tasks like image captioning, visual reasoning, and complex conversations. UniChart Masry et al. (2023), on the other hand, has demonstrated the powerful capabilities of VLMs in the understanding of formatted charts.
With the proliferation of mobile applications, the importance of Graphical User Interfaces (GUIs), which serve as a critical bridge between end users and applications, has increasingly garnered scholarly attention. GUIs, characterized by their structured layouts, rich graphical and textual content, and the inclusion of human operational logic Banerjee et al. (2013), present a complex challenge: Can the success of VLMs be applied to the GUI domain?
Traditional GUI comprehension method focus on conveying the actual user interface interactions and mirror the user’s direct experience with the GUI. For instance, LabelDroid Chen et al. (2020a) utilizes deep learning to predict labels for image-based buttons from a variety of commercial apps available on Google Play. Similarly, TANGO Cooper et al. (2021) employs custom computer vision (CV) and text retrieval techniques to analyze visual and textual information on mobile screens.
However, traditional visual GUI comprehension methods focus mainly on identifying GUI components but fail to fully comprehend the graphical and textual information, layout, and the interaction context within the interface. In contrast, modern approaches such as those employed by ferret-UI and CogAgent, which utilize LVLMs, demonstrate significant advantages in this domain. These methods are designed to process both detailed textual and graphical content, leveraging their pre-trained capabilities. Nevertheless, both ferret-UI You et al. (2024) and CogAgent Hong et al. (2023) frequently encounter limitations by over-relying on their pre-trained knowledge, which leads to a neglect of critical visual content. This often results in the generation of inaccurate or irrelevant responses, indicating a gap in their ability to achieve a balanced and holistic understanding of GUIs Shahgir et al. (2024); Zhang et al. (2023).
@@
  • Performance enhancements on real-world GUI tasks. As shown in experiment  4, we apply our VGA to a GUI comprehension bench, yielding promising results.
@@ 2.2 Large Vision-Language Models
In the field of graphical user interface (GUI) analysis, traditional methods like Ferret-UI You et al. (2024) and CogAgent Hong et al. (2023) predominantly focus on text, largely because their Large Language Model (LLM) components have been thoroughly trained to emphasize textual information. This emphasis can lead to the visual components of the input being overlooked or misinterpreted in the responses, often resulting in the generation of irrelevant answers—a phenomenon known as "hallucination".
@@
Figure 1: Data generation Method
3 Problem in GUI Comprehension
The primary cause of hallucinations in Vision-Language Models (VLMs) can be attributed to their reliance on response patterns learned from traditional Large Language Models (LLMs). In this section, we investigate the occurrence of hallucinations in contemporary GUI models and propose a method to reduce these inaccuracies.
3.1 Hallucination in GUI Comprehension
LLMs are typically trained on massive corpora of pure text data, where they learn to generate responses based on textual context and patterns. When a visual module is introduced, these learned textual response patterns can adversely influence the model’s behavior during multimodal tasks.
Indeed, while Ferrent-UI and CogAgent are adept at understanding specific elements in graphical interfaces, they primarily emphasize textual and positional information when constructing their training datasets. This approach results in limited recognition capabilities for icons and other crucial visual details. Such a bias can lead to misinterpretations or incomplete understanding of GUIs that are visually complex in real-world applications, as reliance on text and layout alone may not adequately capture the full semantics and functionalities of the graphical interfaces. This models can face several types of illusions as follow:
Over-reliance on textual content: LVLMs like Ferrent-UI and CogAgent, heavily trained on text-intensive datasets and focused on text interpretation tasks, tend to overly prioritize textual data when analyzing GUIs. As a result, they might overlook integral visual cues like layout and visual styles, leading to impractical or irrelevant action suggestions within the GUI environment, as shown in Table 1. This text bias affects their capability to deliver effective and holistic interpretations, which is crucial for accurate and user-friendly interface interactions.
Word-to-image coincidences: This issue arises when models draw superficial connections between query words and text appearing on GUI elements, but these elements do not align with the intended functional requirements. As shown in Table 8, if a query includes "start painting," and there are two buttons on the GUI with "start now" and "drawing" in their labels, a model might incorrectly choose a "start now" button when the appropriate action was to "drawing". This misalignment leads to the selection of incorrect actions based on text matches rather than the functional relevance of the elements within the GUI context.
3.2 Referent Method
To address the previously mentioned issue, we propose the referent method. GUI design often uses layout, shape, and color to distinguish elements, providing critical visual details to guide user interactions. By explicitly incorporating coordinates, shapes, colors, and relative positions between elements in the construction of our dataset, we aim to enhance the model’s focus on the image content when generating answers. Thereby reducing the chances of hallucination. During the design of our dataset, we ensured that most responses involving GUI elements include at least one of the following referents, aligning the element information with the image content:
4 GUI Comprehension Dataset
To address the imbalance between image and text in GUI analysis, we propose a high-quality comprehension dataset tailored for LVLM training. This dataset includes detailed annotations of both textual and visual elements, ensuring a balanced and comprehensive training environment.
4.1 Existing General Dataset
Ferret-UI You et al. (2024) and CogAgent Hong et al. (2023), while capable of understanding certain elements in graphical interfaces, primarily focus on extracting and utilizing textual and positional information during the construction of their training datasets. This bias can lead to misunderstandings or inadequate interpretations when dealing with GUIs rich in visual information in actual applications, as relying solely on text and layout might not fully comprehend the semantics and functionalities of the graphical interfaces.
Additionally, since the datasets relied upon by these models are not open to the public, it poses additional challenges for external researchers or developers, as they might struggle to obtain data of similar quality and scale for effective model fine-tuning or further research.
4.2 Data collection
Given that both Ferrent-UI and CogAgent are not open source, the availability of the Rico dataset Deka et al. (2017) represents a valuable resource for the research community involved in mobile app design and development.
The Rico Deka et al. (2017) dataset is a huge GUI dataset created to support research in mobile app design and development, including areas such as GUI design, interaction. It consist of 66k unique GUI screens and 3M elements from 27 categories, over 9.3k applications. Each GUI comprise a screenshot and an augmented Android view hierarchy that capture all of the elements comprising a GUI, their properties, and relationships between them. However, the rich data provided by Rico cannot be directly utilized to train LVLMs as the text data in Rico do not align with human perception in the same manner as VQA datasets. This necessitates the transformation of this data into a format compatible with VQA dataset.
@@ 4.2 Data collection
4.3 Task Design
The dataset includes the following sub-tasks, each complemented by human annotations to elevate the quality and applicability of the training data. Additionally, we have structured the dataset into two distinct categories based on these tasks: the instruction dataset and the conversation dataset. These cater to scenarios of directive compliance and multi-turn dialogue, respectively.
  • Description: This task involves providing a basic description of the GUI’s layout, identifying the function and placement of components. This foundational task facilitates deeper analytical tasks, as all GUI-related tasks rely on a clear understanding of the overall layout.
  • Function: This task requires the model to understand the individual functions of each GUI element, considering not only the overall function but also the relative positions of elements to each other. Identical buttons might serve different purposes depending on their specific context within a GUI. while relative positioning is significant because two elements placed in a specific arrangement might together convey a particular function or information. This enables the model to predict the design intent and functionality of the entire interface, resulting in more contextually aware and precise responses.
4.4 Generation Method
However, the rich data provided by Rico cannot be directly utilized to train LVLMs as the text data in Rico do not align with human perception in the same manner as VQA datasets. This necessitates the transformation of this data into a format compatible with VQA dataset. The Android view hierarchy in the Rico dataset provides detailed descriptions of GUI elements, allowing us to obtain precise positional data and interaction details. Therefore, we use GPT-4 (text-only) to extract this textual information from the Android view hierarchy and GPT-4o to extract visual information from images as shown in Figure 1. Our approaches are as follows:
@@ 4.4 Generation Method
GPT-4 (Text-Only) Generated Tasks: Inspired by the method of LLaVA Liu et al. (2024b), this category consists of dataset generated by GPT-4 (text-only), based on pre-processed Android View Hierarchy data. Leveraging the detailed information available in the Android View Hierarchy, this approach generates a variety of QA pairs that accurately reflect the textual and positional data and interactions.
GPT-4o (Image-Based) Generated Tasks: Inspired by ShareGPT4 Chen et al. (2023), this category comprises dataset generated by GPT-4o (image-based). Given that vision models often struggle with capturing precise positional information of elements, this method focus on relative positions, shapes, and colors of the visual elements.
@@
5 Tuning Script
@@
Foundation Stage:
The Foundation Task dataset is an instruction-following dataset that fix the command and response formats. This stage incorporates direct visual information and employs a fixed format for questions and responses to train the model to correlate with image content. The fixed format is intended to "freeze" the variations in questioning and responding styles, focusing the model’s learning process on understanding and interpreting visual attributes effectively. This controlled environment ensures that the model develops a robust capability to recognize and interpret visual data independently from textual content.
Advanced Comprehensive Stage:
In this stage, we introduce multi-turn dialogues and complex questions from the advanced comprehension dataset. We utilize the Referent Method (detailed in sec 3.2) to directly incorporate intuitive visual information into responses. This method is designed not only to enhance the referencing of visual details within responses but also to align the model’s attention distribution more towards image tokens.It ensures that the model does not merely rely on learned text patterns but actively engages with and interprets the visual context, enabling a deeper understanding and more accurate generation based on the actual GUI layout and visual cues.
@@
Task Progression and Reinforcement:
We organize tasks in a sequence from simple to complex, ensuring that the model progressively builds the necessary foundational skills before training on more advanced tasks. In complex tasks, we start with some related simple tasks to reinforce foundational knowledge, followed by the real response which demands deeper analysis and synthesis of the information gathered from these foundational tasks.
Model First Second Third Average
VGA-no-referent 50.68 48.18 46.14 48.33
VGA-mix-stage 62.05 62.96 62.27 62.42
VGA-only-stage2 67.27 65.00 67.50 66.59
VGA-7b-v1 90.68 90.68 91.17 90.83
Table 3: We conducted ablation experiments on our fine-tuning methods, and the results prove that our fine-tuning methods are effective.
@@ 6.2 Main Result
FAC method analysis: As shown in Figure 2, the orange and purple curves represent the original loss values for models with and without the foundation task stage, respectively. The red and blue curves, fitted with a degree-3 polynomial, illustrate the loss trends. Models trained with the foundation task exhibit more stable convergence and lower loss values. Figure 3 shows that mixing datasets during training leads to unstable convergence. In Table 3, we present the results of ablation experiments on the GUI comprehension bench using the FAC fine-tuning method and those not using the FAC method. The experiment proves that the models using the FAC method have stronger GUI understanding capabilities, consistent with the comparison of the loss curves. In Appendix A.3, we discuses the influence of the foundation stage on model convergence. In Appendix A.4, we conducted ablation studies to evaluate the individual contributions of the two stages to overall model performance. In Appendix A.5, we analyzed the contribution of image tokens when generating answers to demonstrate how FAC enhances the model’s attention to image content.
@@ Acknowledgements
The authors would like to thank the anonymous reviewers for their helpful comments. This research was supported by Bytedance, we would like to extend our gratitude to Bytedance for their assistance and support.